OpenStudy (anonymous):

Input the equation of the given line in standard form. The line with m = 2/3 and passing through (1, 1)

OpenStudy (***[isuru]***):

the standard form y= mx + b m = slope b = intercept u already know the "m" so... y = (2/3) x + b now to find be ... let's use the coordinate (1,1) 1 = (2/3)1 + b b = 1 - (2/3) = 1/3 so... it's DONE! y = (2/3)x + 1/3 hope this will help ya!!